# Compositional Coordination for Multi‑Robot Teams with Large Language Models

### v1

@@ V-B Impact of standardized template
To quantify the impact of using a template to standardize the mission descriptions on LAN2CB, we ablated the template-reformatting stage that aligns user-provided natural language mission descriptions to our predefined mission schema. As reported in Tab. II, this omission results in an average drop of approximately 15% in success rates and causes the error rate to become 2.14 times more frequent across all mission categories. These results highlight the critical role of structured prompting in maintaining robust multi-robot performance.
@@ V-C Real World Demonstrations
We validated our framework through five real-world demonstrations using a team of ground robots across two robot platforms (see Fig. 4).
@@ VI Conclusions
We presented an LLM-based framework, Language to Collective Behavior (LAN2CB), a pipeline that converts natural language into executable Python code for robot teams. Our experiments suggest that LAN2CB enables effective multi-robot coordination and remains flexible and extensible, even for multi-step and long-horizon missions. Additionally, we designed a dataset of natural language mission specifications for multi-robot coordination.
For future work, we aim to improve mission analysis through retrieval-augmented generation (RAG) [50], enabling LLMs to leverage past mission data more effectively. Furthermore, we plan to expand the meta-behavior library to support a broader and more diverse range of capabilities.
VII Acknowledgment
We thank Jiazhen Liu for assistance with the physical deployment.

### v2

@@ V-B Impact of standardized template
To quantify the impact of using a template to standardize the mission descriptions on LAN2CB, we ablated the template-reformatting stage that aligns user-provided natural language mission descriptions to our predefined mission schema. As reported in Tab. II, this omission results in an average drop of $\sim 15\%$ in success rates and causes the error rate to become 2.14×2.14$\times$2.14 × more frequent across all mission categories. These results highlight the critical role of structured prompting in maintaining robust multi-robot performance.
@@ V-C Real World Demonstrations
We validated our framework through four real-world demonstrations [^1] We thank Jiazhen Liu for assistance with the physical deployment. across two robot platforms (see Fig. 4).
@@ VI Conclusions
We presented an LLM-based framework, Language to Collective Behavior (LAN2CB), a pipeline that converts natural language into executable Python code for robot teams. Our experiments suggest that LAN2CB enables effective multi-robot coordination and remains flexible and extensible, even for multi-step and long-horizon missions. Additionally, we designed a dataset of natural language mission descriptions for multi-robot coordination. For future work, we aim to improve mission analysis through retrieval-augmented generation (RAG) [50], enabling LLMs to leverage past mission data more effectively. Furthermore, we plan to expand the meta-behavior library to support a broader and more diverse range of capabilities.